# Split Two-Periodic Aztec Diamond

### v2

@@ Abstract.
Recent advancements have been made to understand the statistics of the Aztec diamond dimer model under general periodic weights [2, CdT24]. In this work we define a model that breaks periodicity in one direction by combining two different two-periodic weightings. We compute the correlation kernel for this Aztec diamond dimer model by extending the methods developed in [3], which utilize the Eynard-Mehta theorem and a Wiener-Hopf factorization. From a form of the correlation kernel that is suitable for asymptotics, we compute the local asymptotics of the model in the different macroscopic regions present. We prove that the local asymptotics of the model agree with the typical two-periodic model in the highest order, however the sub-leading order terms are affected.
@@ 2.1. An Integral Representation of the Correlation Kernel
Figure 6 depicts the contours from the above fromula. The proof of Theorem 2.1 is in Section 3, with some details deferred to Sections 5 and 6.
@@
(17) $\Big{[}\mathbb{K}_{N}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}=-\frac{\mathbb{I}_{m>m^{\prime}}}{2\pi\emph{i}}\oint_{\gamma_{0,1}}\frac{\text{d}z}{z}z^{\xi^{\prime}-\xi}\Phi_{\beta}(z)^{\frac{N}{2}-m^{\prime}}\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\\ +\frac{1}{(2\pi\emph{i})^{2}}\oint_{\gamma_{1}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\beta,1}(w)^{\frac{N}{2}-m^{\prime}}F_{\beta,1}(w)\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\\ +\frac{1}{(2\pi\emph{i})^{2}}\oint_{\gamma_{1}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\beta,2}(w)^{\frac{N}{2}-m^{\prime}}\frac{2}{1+2g_{\alpha,\beta}(w)}F_{\beta,2}(w)F_{\alpha,1}(w)F_{\beta,1}(w)\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\$
@@
This corollary is proven at the end of Section 3. Despite the fact that the initial form of the kernel was more compact, this expression of the kernel is more useful for making comparisons to the typical two-periodic model, which we consider below in Section 2.2. The statement in Corollary 2.1 also allows us to state the kernel succinctly for any $0<m^{\prime}<N$,
@@
(18) $\Big{[}\mathbb{K}_{N}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}=-\frac{\mathbb{I}_{m>m^{\prime}}}{2\pi\textrm{i}}\oint_{\gamma_{0,1}}\frac{\text{d}z}{z}z^{\xi^{\prime}-\xi}\Phi_{\varepsilon(m^{\prime})}(z)^{\frac{N}{2}-m^{\prime}}\Phi_{\varepsilon(m)}(z)^{m-\frac{N}{2}}\\ +\frac{1}{(2\pi\textrm{i})^{2}}\sum_{k=1,2}\oint_{\gamma_{1}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\varepsilon(m^{\prime}),1}(w)^{\frac{N}{2}-m^{\prime}}r_{\varepsilon(m),k}(z)^{m-\frac{N}{2}}F_{\varepsilon(m^{\prime}),1}(w)F_{\varepsilon(m),k}(z)\\ +\frac{1}{(2\pi\textrm{i})^{2}}\sum_{k=1,2}\oint_{\gamma_{1}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\varepsilon(m^{\prime}),2}(w)^{\frac{N}{2}-m^{\prime}}r_{\varepsilon(m),k}(z)^{m-\frac{N}{2}}\\ \times\frac{2}{1+2g_{\alpha,\beta}(w)}F_{\varepsilon(m^{\prime}),2}(w)F_{\overline{\varepsilon}(m^{\prime}),1}(w)F_{\varepsilon(m^{\prime}),1}(w)F_{\varepsilon(m),k}(z)\$
@@
In equation (18) we have also taken the additional step and expanded the matrix $\phi_{\varepsilon}(z)$ in terms of its eigen-decomposition.
@@ 2.2. Comparison to the Two-Periodic Aztec Diamond
(20) $\Big{[}\mathbb{K}_{N}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}=\Big{[}\mathbb{K}_{TP,N}^{\varepsilon}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}\\ +\frac{1}{(2\pi\textrm{i})^{2}}\sum_{k=1,2}\oint_{\gamma_{1}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\varepsilon,2}(w)^{\frac{N}{2}-m^{\prime}}r_{\varepsilon,k}(z)^{m-\frac{N}{2}}\\ \times\frac{2}{1+2g_{\alpha,\beta}(w)}F_{\varepsilon,2}(w)F_{\bar{\varepsilon},1}(w)F_{\varepsilon,1}(w)F_{\varepsilon,k}(z)$
From this expression we see how the correlation kernel of the split model relates to the typically two-periodic model. In order to understand how the local asymptotics of the split two-periodic Aztec diamond differ from the local asymptotics of the typical two-periodic Aztec diamond we only need to understand the asymptotics of the additional two double contour integrals. These will be our main focus when describing the local asymptotics in Section 2.4.
@@ 2.4. Asymptotic Behavior
(25) $\Big{[}\mathbb{K}_{N}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}=\Big{[}\mathbb{K}_{TP,N}^{\varepsilon}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}\\ +\frac{1}{(2\pi\textrm{i})^{2}}\sum_{k=1,2}\oint_{\gamma_{1}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{2}{1+2g_{\alpha,\beta}(w)}F_{\varepsilon,2}(w)F_{\bar{\varepsilon},1}(w)F_{\varepsilon,1}(w)F_{\varepsilon,k}(z)\\ \times\exp\left[N\left(\psi_{2}(w;x,y)-\psi_{k}(z;x,y)\right)+\varphi_{\varepsilon,2}(w;x_{2},y_{2})-\varphi_{\varepsilon,k}(z;x_{1},y_{1})\right]$
@@
(27) $\Big{[}I^{\varepsilon}_{2,k}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}=\frac{1}{(2\pi\textrm{i})^{2}}\oint_{\gamma_{1}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{2}{1+2g_{\alpha,\beta}(w)}F_{\varepsilon,2}(w)F_{\bar{\varepsilon},1}(w)F_{\varepsilon,1}(w)\\ \times F_{\varepsilon,k}(z)\exp\left[N\left(\psi_{2}(w;x,y)-\psi_{k}(z;x,y)\right)+\varphi_{\varepsilon,2}(w;x_{2},y_{2})-\varphi_{\varepsilon,k}(z;x_{1},y_{1})\right]$
@@
(28) $\Big{[}\mathbb{K}_{N}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}=\Big{[}\mathbb{K}_{TP,N}^{\varepsilon}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}\\ +\sum_{k=1,2}\Big{[}I^{\varepsilon}_{2,k}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}$
Thus the local asymptotics of the split model depends on the asymptotics of the typically two-periodic Aztec diamond along with the asymptotics of the contour integrals $I^{\varepsilon}_{2,1}$ and $I^{\varepsilon}_{2,2}$. The local asymptotics of the two-periodic Aztec diamond are well studied, so we will focus on the asymptotic behavior of these two additional contour integrals.
We will utilize the symmetry of the kernel and just consider the asymptotics on the $\alpha$-side of the interface, e.g. when $x<\nicefrac{{1}}{{2}}$. In doing so, we should consider two cases:
@@ 3.1. An Intermediate Correlation Kernel
Where $\gamma_{\text{Br}}$ denotes the contour surround the branch cuts (in the clockwise direction). When we deform the contour $\gamma_{\text{Br}}$ back to $\gamma_{1}$, we once again pick up another single integral term,
(61) $=\oint_{\gamma_{1}}\oint_{\gamma_{0,1}}\frac{\text{d}w\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\beta,2}(w)^{\frac{N}{2}-m^{\prime}}\frac{2F_{\beta,2}(w)F_{\alpha,1}(w)F_{\beta,1}(w)}{1+2g_{\alpha,\beta}(w)}\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\\ -\frac{1}{2\pi\textrm{i}}\oint_{\gamma_{0,1}}\frac{\text{d}z}{z}z^{\xi^{\prime}-\xi}r_{\beta,2}(z)^{\frac{N}{2}-m^{\prime}}\frac{2F_{\beta,2}(z)F_{\alpha,1}(z)F_{\beta,1}(z)}{1+2g_{\alpha,\beta}(z)}\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\\ -\frac{1}{2\pi\textrm{i}}\oint_{\gamma_{0,1}}\frac{\text{d}z}{z}z^{\xi^{\prime}-\xi}r_{\beta,1}(z)^{\frac{N}{2}-m^{\prime}}\frac{2F_{\beta,1}(z)F_{\alpha,2}(z)F_{\beta,2}(z)}{1+2g_{\alpha,\beta}(z)}\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}$
The two single integrals end up canceling out via the same procedure just described (e.g. for one of the integrals, push the contour through the branch cut to the other sheet). Now we have obtained the results presented in Corollary 2.1. ∎

### v3

@@ Abstract.
Recent advancements have been made to understand the statistics of the Aztec diamond dimer model under general periodic weights. In this work we define a model that breaks periodicity in one direction by combining two different two-periodic weightings. We compute the correlation kernel for this Aztec diamond dimer model by extending the methods developed by Berggren and Duits (2019), which utilize the Eynard-Mehta theorem and a Wiener-Hopf factorization. From a form of the correlation kernel that is suitable for asymptotics, we compute the local asymptotics of the model in the different macroscopic regions present. We prove that the local asymptotics of the model agree with the typical two-periodic model in the highest order, however the sub-leading order terms are affected.
@@ 2.1. An Integral Representation of the Correlation Kernel
Figure 6 depicts the contours from the above formula. The proof of Theorem 2.1 is in Section 3, with some details deferred to Sections 5 and 6.
@@
(17) $\Big{[}\mathbb{K}_{N}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}=-\frac{\mathbb{I}_{m>m^{\prime}}}{2\pi\emph{i}}\oint_{\gamma_{0,1}}\frac{\text{d}z}{z}z^{\xi^{\prime}-\xi}\Phi_{\beta}(z)^{\frac{N}{2}-m^{\prime}}\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\\ +\frac{1}{(2\pi\emph{i})^{2}}\oint_{\gamma_{1}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\beta,1}(w)^{\frac{N}{2}-m^{\prime}}F_{\beta,1}(w)\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\\ +\frac{1}{(2\pi\emph{i})^{2}}\oint_{\gamma_{\text{Br}}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\beta,2}(w)^{\frac{N}{2}-m^{\prime}}\frac{2}{1+2g_{\alpha,\beta}(w)}F_{\beta,2}(w)F_{\alpha,1}(w)F_{\beta,1}(w)\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\\ +\frac{1}{2\pi\textrm{i}}\oint_{\gamma_{0,1}}\frac{\text{d}z}{z}z^{\xi^{\prime}-\xi}r_{\beta,1}(z)^{\frac{N}{2}-m^{\prime}}\frac{2F_{\beta,1}(z)F_{\alpha,2}(z)F_{\beta,2}(z)}{1+2g_{\alpha,\beta}(z)}\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}$
We us the notation $\gamma_{\text{Br}}$ to denote a contour surrounding the branch cuts $(-\infty,-\delta^{-2}]\cup[-\delta^{2},0]$ in the clockwise direction, where $\delta=\max\{\alpha,\beta\}$
@@
This corollary is proven at the end of Section 3. Despite the fact that the initial form of the kernel was more compact, this expression of the kernel is more useful for making comparisons to the typical two-periodic model, which we consider below in Section 2.2. As was done in Corollary 2.1, we can modify the second double contour integral in (15) so that the contour $\gamma_{1}$ is deformed to the contour $\gamma_{\text{Br}}$. In doing so, we also get a single contour integral term of a slightly different form,
@@
$\frac{1}{(2\pi\emph{i})^{2}}\oint_{\gamma_{1}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\alpha,2}(w)^{\frac{N}{2}-m^{\prime}}\frac{2F_{\alpha,2}(w)F_{\beta,1}(w)F_{\alpha,1}(w)}{1+2g_{\alpha,\beta}(w)}\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\\ =\frac{1}{(2\pi\emph{i})^{2}}\oint_{\gamma_{\text{Br}}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\alpha,2}(w)^{\frac{N}{2}-m^{\prime}}\frac{2F_{\alpha,2}(w)F_{\beta,1}(w)F_{\alpha,1}(w)}{1+2g_{\alpha,\beta}(w)}\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\\ +\frac{1}{2\pi\textrm{i}}\oint_{\gamma_{0,1}}\frac{\text{d}z}{z}z^{\xi^{\prime}-\xi}r_{\beta,2}(z)^{\frac{N}{2}-m^{\prime}}\frac{2F_{\alpha,2}(z)F_{\beta,1}(z)F_{\alpha,1}(z)}{1+2g_{\alpha,\beta}(z)}\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}$
For brevity, we define
(18) $\Big{[}J(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}=\\ \frac{\mathbb{I}_{m^{\prime}\leq\frac{N}{2}}}{2\pi\textrm{i}}\oint_{\gamma_{0,1}}\frac{\text{d}z}{z}z^{\xi^{\prime}-\xi}r_{\beta,2}(z)^{\frac{N}{2}-m^{\prime}}\frac{2F_{\alpha,2}(z)F_{\beta,1}(z)F_{\alpha,1}(z)}{1+2g_{\alpha,\beta}(z)}\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\\ +\frac{\mathbb{I}_{m^{\prime}>\frac{N}{2}}}{2\pi\textrm{i}}\oint_{\gamma_{0,1}}\frac{\text{d}z}{z}z^{\xi^{\prime}-\xi}r_{\beta,1}(z)^{\frac{N}{2}-m^{\prime}}\frac{2F_{\beta,1}(z)F_{\alpha,2}(z)F_{\beta,2}(z)}{1+2g_{\alpha,\beta}(z)}\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}$
Now we may succinctly state the kernel for any $0<m^{\prime}<N$,
(19) $\Big{[}\mathbb{K}_{N}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}=-\frac{\mathbb{I}_{m>m^{\prime}}}{2\pi\textrm{i}}\oint_{\gamma_{0,1}}\frac{\text{d}z}{z}z^{\xi^{\prime}-\xi}\Phi_{\varepsilon(m^{\prime})}(z)^{\frac{N}{2}-m^{\prime}}\Phi_{\varepsilon(m)}(z)^{m-\frac{N}{2}}\\ +\frac{1}{(2\pi\textrm{i})^{2}}\oint_{\gamma_{1}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\varepsilon(m^{\prime}),1}(w)^{\frac{N}{2}-m^{\prime}}F_{\varepsilon(m^{\prime}),1}(w)\Phi_{\varepsilon(m)}(z)^{m-\frac{N}{2}}\\ +\frac{1}{(2\pi\textrm{i})^{2}}\oint_{\gamma_{\text{Br}}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\varepsilon(m^{\prime}),2}(w)^{\frac{N}{2}-m^{\prime}}\frac{2}{1+2g_{\alpha,\beta}(w)}\\ \times F_{\varepsilon(m^{\prime}),2}(w)F_{\overline{\varepsilon}(m^{\prime}),1}(w)F_{\varepsilon(m^{\prime}),1}(w)\Phi_{\varepsilon(m)}(z)^{m-\frac{N}{2}}\\ +\Big{[}J(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}$
@@ 2.2. Comparison to the Two-Periodic Aztec Diamond
(21) $\Big{[}\mathbb{K}_{N}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}=\\ \Big{[}\mathbb{K}_{TP,N}^{\varepsilon}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}+\Big{[}J(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}\\ +\frac{1}{(2\pi\textrm{i})^{2}}\sum_{k=1,2}\oint_{\gamma_{\text{Br}}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\varepsilon,2}(w)^{\frac{N}{2}-m^{\prime}}r_{\varepsilon,k}(z)^{m-\frac{N}{2}}\\ \times\frac{2}{1+2g_{\alpha,\beta}(w)}F_{\varepsilon,2}(w)F_{\bar{\varepsilon},1}(w)F_{\varepsilon,1}(w)F_{\varepsilon,k}(z)\$
From this expression we see how the correlation kernel of the split model relates to the typically two-periodic model. In order to understand how the local asymptotics of the split two-periodic Aztec diamond differ from the local asymptotics of the typical two-periodic Aztec diamond we only need to understand the asymptotics of the additional two double contour integrals and the single integral term $J(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)$. These will be our main focus when describing the local asymptotics in Section 2.4.
@@ 2.4. Asymptotic Behavior
(26) $\Big{[}\mathbb{K}_{N}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}=\\ \Big{[}\mathbb{K}_{TP,N}^{\varepsilon}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}+\Big{[}J(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}\\ +\frac{1}{(2\pi\textrm{i})^{2}}\sum_{k=1,2}\oint_{\gamma_{\text{Br}}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{2}{1+2g_{\alpha,\beta}(w)}F_{\varepsilon,2}(w)F_{\bar{\varepsilon},1}(w)F_{\varepsilon,1}(w)F_{\varepsilon,k}(z)\\ \times\exp\left[N\left(\psi_{2}(w;x,y)-\psi_{k}(z;x,y)\right)+\varphi_{\varepsilon,2}(w;x_{2},y_{2})-\varphi_{\varepsilon,k}(z;x_{1},y_{1})\right]$
@@
(28) $\Big{[}I^{\varepsilon}_{2,k}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}=\frac{1}{(2\pi\textrm{i})^{2}}\oint_{\gamma_{\text{Br}}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{2}{1+2g_{\alpha,\beta}(w)}F_{\varepsilon,2}(w)F_{\bar{\varepsilon},1}(w)F_{\varepsilon,1}(w)\\ \times F_{\varepsilon,k}(z)\exp\left[N\left(\psi_{2}(w;x,y)-\psi_{k}(z;x,y)\right)+\varphi_{\varepsilon,2}(w;x_{2},y_{2})-\varphi_{\varepsilon,k}(z;x_{1},y_{1})\right]$
@@
(29) $\Big{[}\mathbb{K}_{N}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}=\Big{[}\mathbb{K}_{TP,N}^{\varepsilon}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}\\ \Big{[}J(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}+\sum_{k=1,2}\Big{[}I^{\varepsilon}_{2,k}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}$
Thus the local asymptotics of the split model depends on the asymptotics of the typically two-periodic Aztec diamond along with the asymptotics of the contour integrals $J$, $I^{\varepsilon}_{2,1}$ and $I^{\varepsilon}_{2,2}$. We will start with a statement about $J$.
Proposition 2.1.
For $(m^{\prime},\xi^{\prime})$ and $(m,\xi)$ defined by equations (24) and (25), we have
(30) $|J(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)|\leq c_{2}e^{-c_{1}N}$
where $c_{1}$ and $c_{2}$ are positive constants.
We will utilize the symmetry of the double contour integrals and just consider the asymptotics on the $\alpha$-side of the interface, e.g. when $x<\nicefrac{{1}}{{2}}$. In doing so, we should consider two cases:
@@ 3.1. An Intermediate Correlation Kernel
Where $\gamma_{\text{Br}}$ denotes the contour surround the branch cuts (in the clockwise direction). ∎